Find the magnitude and directions of the current

쇼 15V 50L
Answer #1

let the current be I1 in top loop and be i2 in bottom loop

writing the loop equation we have

-9I1 - 6(I1 -I2) + 15 = 0 ---(1)

-15 -6(I2- i1) - 5I2 -12 = 0 ----(2)

ADDING Both equation we have

9I1 + 5I2 = -12  

I1 = (-12 - 5I2) / 9

-15*-(12 + 5I2) / 9 + 6I2 +15 = 0

35 + 14.33I2 = 0

I2 = -2.44 A

I1 = -(15 + 5*-2.44) / 9

= -0.31 A